OpenStudy (anonymous):

Find the solution to the differential equation that satisfies the initial conditions y(0)=50

OpenStudy (usukidoll):

where is the differential equation?

OpenStudy (anonymous):

dy/dt=ky

OpenStudy (anonymous):

k is a constant

OpenStudy (usukidoll):

hmmm separable equation?

OpenStudy (usukidoll):

1/y dy = k dt... ln y = k t y = e^kt ?

OpenStudy (anonymous):

how do you include the initial condition?

OpenStudy (anonymous):

it's been a year since my calculus ab course and now i'm in physics c :/

OpenStudy (usukidoll):

Y(0) = 50 soo the ?? 50 = e^k(0) ???

OpenStudy (usukidoll):

that doesn't look right ... that's 1 = 50 nooooo

OpenStudy (usukidoll):

ln y = k t

OpenStudy (usukidoll):

e^lny = e^kt djfslakjdaljfd

OpenStudy (anonymous):

would it be y=50e^k*t then 100=50e^k*5 so k=0.1386 y=50e^90.1386)8t??

OpenStudy (anonymous):

that 8 should be *

OpenStudy (anonymous):

y+50e^(0.1386)*t